### v4

@@ 5.5 Visualization
Focus Weight. Fig. 7 depicts the focus weights of selected questions. Notably, questions seeking information about specific actions (e.g., “win a point”) heavily rely on short-term information (focus length 3). In contrast, questions like “How many times does the right team do ‘block’?” necessitate a stronger emphasis on long-term information. Questions pertaining to the relationships between two actions show a preference for mid-term information, offering insights into the cause and effect of actions. Intuitively, the mid-term focus length can provide more nuanced information on the interactions between actions. This visualization affirms the effectiveness of our model, which dynamically integrates different focus scales based on the diverse temporal dependencies required by various questions.
@@ 7 Limitation and Future Work
In our study, we introduce the innovative Auto-Focus Transformer (AFT) designed to capture intricate dependencies across frames within diverse temporal windows, catering to a range of question types. Unlike the conventional Transformer model, the AFT demands increased computational resources, scaling linearly with the number of focuses. Although we have optimized the implementation for high parallelization, it may pose challenges on devices with constrained memory capacities. Looking ahead, our research will focus on devising a novel architectural framework that retains the distinctive features of AFT while ensuring computational efficiency, addressing the limitations associated with memory constraints.
@@ 8 Ethical Consideration
In developing our Sports-QA dataset, sourced from MultiSports and FineGym, we prioritized ethical considerations, adhering to data privacy and usage standards while ensuring compliance with the terms of the original datasets. To address potential biases, we actively promoted diversity in sports categories and specific sports actions, aiming for a balanced representation. Mindful of the environmental impact, we optimized our data processing to minimize resource consumption. Acknowledging the potential for misuse, we emphasize the dataset’s intended use in positive sports analytics applications, such as performance analysis. Our commitment to transparency and reproducibility is evident in our methodology and dataset accessibility, ensuring that our contribution to sports analytics and computer vision is both responsible and beneficial.

### v5

@@ 5.5 Visualization
Focus Weight. Fig. 7 depicts the focus weights of selected questions. Notably, questions seeking information about specific actions (e.g., “win a point”) heavily rely on short-term information (foccal length 3). In contrast, questions like “How many times does the right team do ‘block’?” necessitate a stronger emphasis on long-term information. Questions pertaining to the relationships between two actions show a preference for mid-term information, offering insights into the cause and effect of actions. Intuitively, the mid-term focal length can provide more nuanced information on the interactions between actions. This visualization affirms the effectiveness of our model, which dynamically integrates different focus scales based on the diverse temporal dependencies required by various questions.
@@ 7 Limitation and Future Work
Our proposed model does exhibit zero-shot generalization ability, precisely because it is not pretrained on large-scale multimodal datasets—this absence of prior exposure to such data is what underscores its capacity to generalize to unseen scenarios without explicit training. That said, we recognize the value in further exploring the nuances of this generalization capability, and we plan to delve deeper into this aspect as part of our future work to provide more comprehensive insights. Additionally, we convert answers into 191 categories in our dataset, which limits free-form generation and semantic diversity. This formulation merely serves as a practical intermediate step between rigid classification and fully free-form generation. We plan to explore evaluation methods that capture open-ended responses, such as semantic similarity metrics or hybrid protocols combining categorical and free-form assessments.
@@ 8 Ethical Consideration
In developing our Sports-QA dataset, sourced from MultiSports and FineGym, we prioritized ethical considerations, adhering to data privacy and usage standards while ensuring compliance with the terms of the original datasets. To address potential biases, we actively promoted diversity in sports categories and specific sports actions, aiming for a balanced representation. Mindful of the environmental impact, we optimized our data processing to minimize resource consumption. Acknowledging the potential for misuse, we emphasize the dataset’s intended use in positive sports analytics applications. Our commitment to transparency and reproducibility is evident in our methodology and dataset accessibility, ensuring that our contribution to sports analytics and computer vision is both responsible and beneficial.
@@
Acknowledgments
This research was particularly supported by the Australian Government through the Australian Research Council’s DECRA funding scheme (Grant No.: DE250100030), Discovery Project funding scheme (Grant No.: DP260100218, DP260101891), the National Natural Science Foundation of China (No. 62372491), and the Guangdong Basic and Applied Basic Research Foundation (2022B1515020103).